1. The physician’s order reads: 1.8 L of NS to infuse over 15 hours
a. How many mL/ h will the IV need to run at in order to deliver the solution as ordered?

2. The physician’s order reads: Give antibiotic in 30 mL D5W over 15 minutes
a. How many mL/ h will the IV need to run at in order to deliver the solution as ordered?

3. 1200 mL D5W IV is ordered to infuse in 10 hours by infusion pump. Calculate the flow rate in milliliters per hour.

4. 600 mL of antibiotic is to be infused over the 180 minutes by an infusion pump. Calculate the flow rate (mL per hour).

5. The physician’s order reads: Dopamine 20 mcg/Kg/minute. The bag is labeled Dopamine 100 mg/50 ml. The patient weighs 88 lbs. How fast will the IV run in mL/hr to give the dose?

read each questions thoroughly and answer correctly..Everything is straightforward and clear. nothing isunclear
Answer #1

ANSWER.

1) Calculation of ml/hr.

          ml/hr = Volume(ml) / Time(hr).

   Here

      Volume = 1.8 L.

Convert L to ml for calculation

          1 L = 1000 ml.

So

    Volume(ml) = 1.8 * 1000

                        = 1800 ml.

        Time(hr) = 15 hr.
As per above formula

          ml/hr = 1800/15

                   = 120 ml/hr.

2) Calculation of ml/hr.

          ml/hr = Volume(ml) / Time(hr).

   Here

   Volume(ml) = 30 ml.

   Time = 15 min

Convert 15 min to hr for calculation

                60 min = 1 hr

                15 min = X hr

         X * 60 = 15 * 1

              X = 15/60

                 = 0.25 hr.

   So

         Time(min) = 0.25 hr.

As per above formula

             ml/hr = 30/0.25

                      = 120 ml/hr.

3) Calculation of ml/hr.

           ml/hr = Volume(ml) / Time(hr).

     Here

          Volume(ml) = 1200 ml.

          Time(hr) = 10 hr.

As per above formula

        ml/hr = 1200/10

                 = 120 ml/hr.

4) Calculation of ml/hr.

            ml/hr = Volume(ml) / Time(hr).

       Here

           Volume(ml) = 600 ml.

           Time(hr) = 3 hr (180 min).

As per above formula

          ml/hr = 600 / 3

                   = 200 ml/hr.

5) Calculation of ml/hr.

            ml/hr = Volume(ml) / Time(hr).

    Calculation of volume

        Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q)

       Here

                 Ordered dose = 20 mcg/kg/min.

                 Weight of the patient = 88 lbs.

   Convert lbs to kg for calculation

               1 lbs = 0.454 kg

      So

            Weight of the patient in kg = 88 * 0.454

                                                       = 39.9 kg.

            Desired dose(D) for this patient = 20 * 39.9

                                                                = 798 mg.

                      Dose in hand(H) = 100 mg.

      Convert mg to mcg for calculation

                       1 mg = 1000 mcg

          So

                Dose in hand(H) = 100 * 1000

                                            = 100000 mcg.

                            Quantity(Q) = 50 ml.

    As per above formula

                Volume(ml) of drug required = D/H * Q

                                                              = 798/100000 * 50

                                                              = 0.00798 * 50

                                                              = 0.4 ml

                              Time(min) = 1 min.

     Convert ml/min to ml/hr by multiply with 60

                         ml/hr = 0.4/1 * 60

                                  = 24 ml / hr.